OpenStudy (anonymous):

-5/4 x - 3/5y= -1/6 15/16x 9/20y= 1/8 one solution, no solution, infinite number of solutions???

OpenStudy (anonymous):

Is the first equation this one \[-\frac{5}{4 x}-\frac{3}{5 y}=-\frac{1}{6} \] ? Can you be more precise about the second equation by using parentheses?

jhonyy9 (jhonyy9):

the second equation is very difficile

OpenStudy (anonymous):

Is there a sign between 15/16x and 9/20y ?